OpenStudy (anonymous):

Simplify the given expression (1 + 2i) • (5 − 3i) ?

OpenStudy (anonymous):

5 - 3i + 10i - 6i 5 + i

OpenStudy (anonymous):

simplifying gives 5-3i+10i+6 which is =11+7i

OpenStudy (anonymous):

how'd you get 11 out of 5 being the only coefficient?

OpenStudy (anonymous):

2i * -3i=-6(-1)=6 since i^2 is -1

OpenStudy (anonymous):

so, (1+2i)(5-3i)=5-3i+10i+6 where 6 is goten as above

OpenStudy (anonymous):

this isn't an equation though, it's an expression, therefore cant be turned into the a, b, c.

OpenStudy (anonymous):

sorry change = to 

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you change equal to sign to an implication sign if you wish

OpenStudy (anonymous):

but the concept remains the same